### 29 Dec 2023 v2

# Through the eyes of a reader and science communicator: science in the mainstream and in the genre literature of yesterday and today

Abstract
For most writers the science is either an exotic setting or a source of thrilling conflict that would drive the story forward. For a communicator it is the other way around – the science is neatly wrapped in a package of literary tools that make it “invisible” while it remains tangible and most importantly – it can be conveyed to the reader in understandable terms. There are many examples showing how these seemingly contradicting goals can complement each other successfully. I will review how the science was communicated by mainstream and genre writers of yesterday and today, and in different (not necessarily anglophone) cultures. I will bring forward the best and the worst examples that illuminate various astronomical concepts. Finally, I will discuss how we can use them both in outreach and education. Contrary to many similar summaries I will concentrate on some often overlooked mainstream literary examples, including the plays The Physisists by Friedrich Dürrenmatt and Copenhagen by Michael Frayn, the novel White Garments by Vl. Dudintsev and even an episode of the Inspector Morse TV show, featuring scientists. I will also mention in passing a few less well known genre books.
1 Introduction
The public image of sciences and scientists is relevant, especially for the fundamental science that depend on public support and funding. However, the sciences are often subject to heavy stereotyping and adversity and the improvement of public perception can be a daunting task. The problem is recognized: [2] polled the American public about impression of people according to their jobs, discriminating between “apparent intent (warmth) and capability (competence).” The outcome places researchers at the low-warmth/high-competence end, together with accountants, lawyers, engineers and chief executive officers. Without underestimating the importance of these professions, we can safely assume that at least some of them need public image improvement. Turning to the historic science figures, [1] analyzed images of Greek scientists to trace the evolution of how they are perceived by the public, and found that while the public image of the Greek scientists have improved, it is till dominated by the traditional stereotypes. Studies of the public image of scientists and its implications for teaching of sciences in various cultures are also available, e.g., [5].
The depiction of sciences and scientists in arts deserves a special attention because of the powerful emotional influence that arts can exert on the public [6, 7]. Here I review some popular works of art that describe how science and scientists work, and a few less well known examples from the Eastern European add in particular from the Bulgarian cultural lore.
2 Scientists though the eye of the artists
Scientists are often featured in popular arts. To remind a few, almost household names: Susan Calvin, Indiana Johns, Victor Frankenstein, Dr. Strangelove, Capt. Nemo, Emmett Brown, Eleanor Arroway, Spock, Dr. No, Sheldon Cooper, Dana Scully, Dr. Kelvin, and even Sherlock Holmes. I have intentionally included in the the list both males and females, protagonists and antagonists, to demonstrate the diversity of the characters, the variety of their motives and the range of their actions. A lesser known name is that of Dr. Kelvin, the psychologist from the novel Solaris by Stanislaw Lem (and two motion pictures based on this book).
What brings these characters together is the analytic approach to the problems they face and the application of the scientific method; what divides them is that some use the scientific tools to expand our knowledge and to make the people’s lives easier and others – to harm people and society.
For writers and screenwriters invoking scientists is often an easy way to create all-powerful menacing figures that would drive the story forward without the need to invest much though and imagination. Further fuel into this negative outlook to science is provided by two notable novels addressing the devastating consequences that the lack of scientific integrity can bring – Solar by Ian McEwan and The White Garments by Vladimir Dudintsev. The former has a modern setting and follows the life story of an eco-scientist who makes some very questionable decisions. While at its core this is a novel about integrity, the topic is shown to exceeds the limits of the guarded and somewhat confined scientific environment, it is shown as a problem of world shattering proportions. The second book appeared in the first years of the Soviet Perestroyka, when Eastern Europe opened up and faced many issues, including to its own past. It describes a conflict dating from the early post-WWII years while Stalin was still in power and the science succumbing to or rather was trying to fight a loosing battle against ideology. The novel asked the uncomfortable question what can one do, if the research suggest that the world pretty much disagrees with the established ideology and with the party line. Silence and internal immigration is one answer. Summarizing, these two books offer less than favorable images of scientists, and they are a bit too big of a reading to be useful topic starters that would introduce students to ecology and biology,
Another example of a fictionalized presentation of scientists in a completely different environment comes from the Bulgarian writer Dimitar Peev (1919-1996) who published in 1966 a space faring novel The Photon Starship. The action is set in a communist Utopia and follows a sub-light speed trip to the nearest star system Alpha Cen. Surprisingly for a books of that era it contains very little ideology and a lot of scientific speculations, some – quite relevant (for how these can be used as educational topic raisers see [3]). The entire expedition consists of highly motivated scientists, so many that nobody really stands out. This was a concession to the politics back them – the individual did not matter, the collective and the society ruled. However, the increasing average number of authors per paper today may be returning us back to this idea. We can safely say that Peev has adopted a science-optimistic stand. His book offers insights into planet formation and stellar binary formation, although critical comparison with the modern understanding of these topics is needed.
In Stars and Waves Roberto Maiolino, a professional astronomer himself, offers an insider’s look at the life in academic environment, a view of the intricate research process, the motivation of the investigators and even some enticing description of the technology and methodology, both in the present day world leading astronomical observatory, and in big observatory behind the Iron Curtain from a few decades in the past. The novel takes the readers on a lavishly illustrated trip across the largest modern astronomical facilities and research centers. It starts like a mystery/investigation procedural. However, it soon enters the territory of international politics with hints of espionage, with detours towards personality studies that include big egos, hubris and un-moderated ambitions. Gradually some science-fictional elements build up, and at the end we find ourselves with a full fledged scifi genre work. The author does not shy away from some painful issues that still – sadly – plague the modern sciences, including the acceptance (or lack thereof) of women and ending with the outright stealing of other people’s ideas.
However, the novel should be taken with a big grain of salt as a guide to real-life university life, because the author has made some big concessions to build suspense and boost conflicts. Without revealing too much content, I can only tell that Laura, the graduate student protagonist, suffers from stage fright and have difficulties giving presentations. Every institution that I have studies or worked at in my career had a student seminar, where the students would present their work to each other, in a friendly environment, just to overcome this particular problem. Still, Stars and Waves can be a good vehicle to introduce the audience to academic life, provided such points are addressed and discussed. Unlike the books mentioned earlier it offers much more balanced view. The novel is an excellent introduction into a discussion of a wide range of astronomy and astrobiology topics – from astroclimate and telescope design to the searches for bio- and technosignatures.
The most popular TV series has been less than merciful to the scientists they have portrayed. Inspector Morse and it sequel Lewis turn Oxford into a dangerous place with a high crime rate. The academics are frequently involved, often shown in negative light, being involved in gambling, extramarital affairs and now and then – murders. The protagonists often make a point of the hubris that the scientists tend to exhibit both in their work and personal lives (even though detective himself is not completely alien to this character trait). The Big Bang Theory makes fun of the scientist – without having to apply much imagination to the script, because the real life suffices to make the audience laugh. At best the show evokes sympathy to the “poor” scientists who are portrayed unable to coop with the challenges of everyday life. Both these TV shows paint negative portraits of scientists, albeit from very different angles. Inspector Morse can hardly help as a science intro, but The Big Bang Theory often touches various science questions in detail.
The theatrical stage has brought is two notable examples of science portrayals – the plays Copenhagen and The Physicists. The first, written by Michael Frayn, is based on a real-life visit that Werner Heisenberg paid to Niels Bohr and his wife Margrethe Bohr. There are differences in the accounts the participants gave afterwords. Clearly, they discussed the atomic bomb, the implications and the ethics of building it. After the meeting Heisenberg went on to work on the Nazi nuclear program, together with Otto Hahn, Carl Friedrich von Weizsäcker and other prominent physicists. Bohr escaped to UK and later to US. Overall, this is a thoughtful piece that puts forward a picture of the scientific word that is aware of the consequences that the researchers will have to face in their work [^1] The movie Oppenheimer that premiered after the EAS 2023 addresses similar issues and it probably helped the public to understand and sympathize with the scientists more than any of the works mentioned here.. While Copenhagen does touch upon the basic nuclear physics, it is of more help for discussions on scientific ethics.
The Dürrenmatt play is different. It starts as a farce, with patients in a psychiatric yard that imagine themselves as Newton and Einstein, but true to the author’s love of surprises, they turn out to be something different yet. The action resonates with the Cold War situation of the day and it is not really a statement about science and scientists, making it less relevant for our discussion.
Finally, I would like to mention four relatively recent collections that offer a look at the science and scientists from insiders: Diamonds in the Sky (ed. Mike Brotherton; 2009), A Kepler’s Dozen: Thirteen Stories About Distant Worlds That Really Exist (eds. Steve B. Howell & David Lee Summers; 2013), Science Fiction by Scientists (ed. Mike Brotherton; 2017), and Life Beyond Us (eds. Julie Nováková, Lucas K. Law & Susan Forest; 2023). Not surprisingly, the authors paint researchers in most positive light, as ethical problem solvers. Almost all of these were designed to help in classroom: the stories are short and selected to capture the attention of students; in some cases the literary pieces are accompanied with non-fictional essays that explain in popular terms the science that the stories touch upon.
3 Discussion and conclusions
Our survey how research scientist are drawn in arts is broad, coarse and certainly not pretending for completeness in any meaningful way. Unfortunately, even this quick overview drives us towards a rather grim conclusion: with a few rare exceptions the researchers are seen either as logical but cold and hubristic beings devoid of human qualities, who can not be relied upon, or as ill-adjusted rueful loners. The power of fiction is to influence people, either on conscious or on subconscious level, and here it works toward undermining the public trust in scholarly figures. The exceptions almost exclusively come from the pen of a scholarly figures and reach limited audience.
Intensive outreach activities are the obvious answer to this problem (e.g., [4]). Usually, the outreach aims at spreading the scientific knowledge among the general public, but others and no less important goals are creating more science friendly environment and putting forward a positive image of the scientists. The outreach should bringing out the idea that doing science is an admirable and accessible human activity, deserving respect.
Acknowledgements
This is an extended write up of a contributed talk presented at the European Astronomical Society (EAS), Special Session 39 “Sci-Art: Communicating Science Through Art”, held held in Kraków, Poland on Jul 10-14, 2023.

### 28 Apr 2024 v3

@@
 Finally, I will discuss how we can use them both in outreach and education.
-Contrary to many similar summaries I will concentrate on some often overlooked mainstream literary examples, including the plays The Physisists by Friedrich Dürrenmatt and Copenhagen by Michael Frayn, the novel White Garments by Vl.
+Contrary to many similar summaries I will concentrate on some often overlooked mainstream literary examples, including the plays The Physicists by Friedrich Dürrenmatt and Copenhagen by Michael Frayn, the novel White Garments by Vl.
 Dudintsev and even an episode of the Inspector Morse TV show, featuring scientists.
@@
 Here I review some popular works of art that describe how science and scientists work, and a few less well known examples from the Eastern European add in particular from the Bulgarian cultural lore.
-2 Scientists though the eye of the artists
+2 Scientists through the eye of the artists
 Scientists are often featured in popular arts.
-To remind a few, almost household names: Susan Calvin, Indiana Johns, Victor Frankenstein, Dr.
+To remind a few, almost household names: Susan Calvin, Indiana Jones, Victor Frankenstein, Dr.
 Strangelove, Capt.
@@
 Kelvin, and even Sherlock Holmes.
-I have intentionally included in the the list both males and females, protagonists and antagonists, to demonstrate the diversity of the characters, the variety of their motives and the range of their actions.
+I have intentionally included in the list both males and females, protagonists and antagonists, to demonstrate the diversity of the characters, the variety of their motives and the range of their actions.
 A lesser known name is that of Dr.
@@
 The novel asked the uncomfortable question what can one do, if the research suggest that the world pretty much disagrees with the established ideology and with the party line.
-Silence and internal immigration is one answer.
-Summarizing, these two books offer less than favorable images of scientists, and they are a bit too big of a reading to be useful topic starters that would introduce students to ecology and biology,
+Silence and “internal immigration” [^1] A term used in the former Eastern Europe to signify a person’s refusal to participate in public life; a refusal to use social media is a modern day analog of this behaviour.
+is one answer.
+Summarizing, these two books offer less than favorable images of scientists, and they are a bit too big of a reading to be useful topic starters that would introduce students to ecology and biology.
 Another example of a fictionalized presentation of scientists in a completely different environment comes from the Bulgarian writer Dimitar Peev (1919-1996) who published in 1966 a space faring novel The Photon Starship.
 The action is set in a communist Utopia and follows a sub-light speed trip to the nearest star system Alpha Cen.
-Surprisingly for a books of that era it contains very little ideology and a lot of scientific speculations, some – quite relevant (for how these can be used as educational topic raisers see [3]).
+Surprisingly for a book of that era it contains very little ideology and a lot of scientific speculations, some – quite relevant (for how these can be used as educational topic raisers see [3]).
 The entire expedition consists of highly motivated scientists, so many that nobody really stands out.
@@
 Without revealing too much content, I can only tell that Laura, the graduate student protagonist, suffers from stage fright and have difficulties giving presentations.
-Every institution that I have studies or worked at in my career had a student seminar, where the students would present their work to each other, in a friendly environment, just to overcome this particular problem.
+Every institution that I have studied or worked at in my career had a student seminar, where the students would present their work to each other, in a friendly environment, just to overcome this particular problem.
 Still, Stars and Waves can be a good vehicle to introduce the audience to academic life, provided such points are addressed and discussed.
-Unlike the books mentioned earlier it offers much more balanced view.
+Unlike the books mentioned earlier it offers a much more balanced view.
 The novel is an excellent introduction into a discussion of a wide range of astronomy and astrobiology topics – from astroclimate and telescope design to the searches for bio- and technosignatures.
@@
 The academics are frequently involved, often shown in negative light, being involved in gambling, extramarital affairs and now and then – murders.
-The protagonists often make a point of the hubris that the scientists tend to exhibit both in their work and personal lives (even though detective himself is not completely alien to this character trait).
+The protagonists often make a point of the hubris that the scientists tend to exhibit both in their work and personal lives (even though the detective himself is not completely alien to this character trait).
 The Big Bang Theory makes fun of the scientist – without having to apply much imagination to the script, because the real life suffices to make the audience laugh.
@@
 Inspector Morse can hardly help as a science intro, but The Big Bang Theory often touches various science questions in detail.
-The theatrical stage has brought is two notable examples of science portrayals – the plays Copenhagen and The Physicists.
+The theatrical stage has brought us two notable examples of science portrayals – the plays Copenhagen and The Physicists.
 The first, written by Michael Frayn, is based on a real-life visit that Werner Heisenberg paid to Niels Bohr and his wife Margrethe Bohr.
@@
 After the meeting Heisenberg went on to work on the Nazi nuclear program, together with Otto Hahn, Carl Friedrich von Weizsäcker and other prominent physicists.
-Bohr escaped to UK and later to US.
+Bohr escaped to the UK and later to the US.
 Overall, this is a thoughtful piece that puts forward a picture of the scientific word that is aware of the consequences that the researchers will have to face in their work [^2] The movie Oppenheimer that premiered after the EAS 2023 addresses similar issues and it probably helped the public to understand and sympathize with the scientists more than any of the works mentioned here..
@@
 3 Discussion and conclusions
-Our survey how research scientist are drawn in arts is broad, coarse and certainly not pretending for completeness in any meaningful way.
+This survey how research scientist are drawn in arts is broad, it lacks resolution and certainly does not claim completeness in any meaningful way.
 Unfortunately, even this quick overview drives us towards a rather grim conclusion: with a few rare exceptions the researchers are seen either as logical but cold and hubristic beings devoid of human qualities, who can not be relied upon, or as ill-adjusted rueful loners.
@@
 Acknowledgements
-This is an extended write up of a contributed talk presented at the European Astronomical Society (EAS), Special Session 39 “Sci-Art: Communicating Science Through Art”, held held in Kraków, Poland on Jul 10-14, 2023.
+This is an extended write up of a contributed talk presented at the European Astronomical Society (EAS), Special Session 39 “Sci-Art: Communicating Science Through Art”, held in Kraków, Poland on July 10-14, 2023.